Assume that the population of heights of female college students is approximately normally distributed with mean m of 66 inches and standard deviation s of 4.00 inches. Show all work.
(A) Find the proportion of female college students whose height is greater than 65 inches.
(B) Find the proportion of female college students whose height is no more than 65 inches
*Where do I begin. I have been working on this problem since 5 and can't figure it out. PLEASE HELP :O(

Answer by stanbon(75887)

Assume that the population of heights of female college students is approximately normally distributed with mean m of 66 inches and standard deviation s of 4.00 inches. Show all work.
---
Comment: You measure table length with a tape measure. You measure
probabilities with a z-scale.
------
(A) Find the proportion of female college students whose height is greater than 65 inches.
z(65) = (65-66)/4 = -0.2500
P(x > 65) = P(z > -0.2500) = normalcdf(-0.25,100) = 0.5987
----------------------------------
(B) Find the proportion of female college students whose height is no more than 65 inches
This is the complement of Problem "A":
P(x <= 65) = 1-0.5987 = 0.4013
=================
Cheers,
Stan H.
===============